Dublin University Rifle Club

Version Date : April 23, 2002 Filename : Status :

Dublin University Rifle Club Match Secretary’s Report

Mark Dennehy Match Secretary

April 23, 2002

1 Attendance Summary

This year, 24 external competitions were attended nationally, four at international level (including two World Cups) and nine internal competitions. Attendance at all of these (except for the Internationals, by virtue of the standard required) was exceptionally high - often swamping the hosting club with scoring and administrative work! This level of attendance does not go unnoticed at any level in the sport and is a very positive public image for the Club and College - those of you who participated in these competitions have just reason to feel proud.

2 DURC competitions hosted

DURC has only hosted one Open competition this year so far, due to the Foot & Mouth crisis which has decimated the outdoor shooting calenday this year. Our Indoor Air Rifle Open was a resounding success however - and after the debacle of the previous Open, it is a welcome sight to see that our efforts were not unnoticed - we have in fact, been requested to do the logistics and administration for the National Air Rifle Championships this weekend. A lot of thanks have to go to Keith White for his computer program and his seven-hour scoring stint at the open for this, as well as to the range officers who assisted during the day - your efforts were noticed by one and all!

The Free Rifle Open is slated to take place this year, just as soon as the Fassaroe range is reopened.

3 External Competitions

3 External Competitions

3.1 National Level

3.1.1 Wicklow 50m Open BRC, May 7, 2000

Name  Class Final Score  Notes
Ray Reilly  A  579  7^{th} Overall
Mark Dennehy C  545  3^{rd} Class C, 23^{rd} Overall

3.1.2 Fassaroe Shooting Club (F.S.C.) 50m Open FSC, June 17/18, 2000 50m Prone Rifle :

Name  Class Final Score  Notes
Ray Reilly  A  577  5^{th} Overall
Ian Marnane C  565  1^{st} Class C, 10^{th} Overall
Mark Dennehy C  554
John Keeney C  552

50m 3P Rifle:

Name  Class Final Score  Notes
Mark Dennehy D  370/176/307 : 853  3rd Place

(out of 3 shooters!)

100yd Pairs :

Name  Individual Score Final Score  Notes
Ray Reilly  367
Alan Smyth 366 733 5^{th} Place
Mark Dennehy 377
John Keeney 372  749  3^{rd} Place

3.1.3 National 50m Championships FSC July 1/2

Name  Class Final Score  Notes
Ray Reilly  A  584 (686.1 Final) National Champion!
Mark Dennehy C  551  2^{nd} Class C
John Keeney C  552  3^{rd} Class C
Shane Kelly  D  431  3^{rd} Class D

3.1.4 Comber Open

Comber Rifle Club August 5/6, 2000

50m Prone Rifle :

Name  Class Score Notes
Ray Reilly  A  578
Mark Dennehy C  550
John Keeney C  527

100yd Pairs :

Name  Class Score  Notes
Ray Reilly  A  377
Mark Dennehy C  381  1^{st}Overall100yds
John Keeney C  346

50m Teams:

Name  Score Notes
Ray Reilly  579
Mark Dennehy 555
John Keeney  582
Patrick O’Brien 549  Honorary Member!

Sadly, last place!

3.1.5 Leinster Open

RRPC Sept 3, 2000 25yd Prone Rifle

Name  Class Final Score  Notes
Ray Reilly  A  585  2^{nd} Class A & Overall
Mark Dennehy  D  576  1^{st} Class D! & 6^{th} Overall
Deborah Mehigan D  477
Michael Madigan D  462

3.1.6 EARC Festival

EARC Oct 7/8, 2000

Air Rifle:

Name  Class Final Score  Notes
Mark Dennehy C  514  1^{st} Class C
John Keeney D  476
Alan Smyth  C  471

50m Rifle Prone:

Name  Class Final Score Notes
Ray Reilly A  582

Confined Aggregate

Name  Final Score Notes
Mark Dennehy 1270  4^{th} Place

3.1.7 UCD Open 1

UCDRC Oct 28/29, 2000 10m Air Rifle

Name  Class Final Score  Notes
Mark Dennehy  C  505  1^{st} Class C
Alan Smyth  C  454  2^{nd} Class C
Aisling Broderick D  501  2^{nd} Class D
Jason Sibley  D  451  3^{rd} Class D
Colm Casserlly  D  387

3.1.8 WAIDAM Open

UCDRC Nov 11/12, 2000 10m Air Rifle

Name  Class Final Score  Notes
Alan Smyth  C  482  1^{st} Class C
Mark Dennehy C  481  2^{nd} Class C
John Keeney D  457
Jason Sibley  D  438
Shane Kelly  D  406

3.1.9 DURC Air Rifle Open

UCDRC Nov 26, 2000 10m Air Rifle

Name  Class Final Score  Notes
Mark Dennehy  C  517(602.9)  1^{st} Class C
Aisling Broderick C  503  3^{rd} Class C
Steven Watterson A  562  Enough said ...
Alan Smyth  C  472
Jason Sibley  D  470  2^{nd} Class D
Kenneth Suzuki D  439
Ben Jones  D  428
Shane Kelly  D  405

3.1.10 Turkey Shoot

RRPC Dec 3, 2000 Invitational Shoot 25yd Prone Rifle

Name  GunScore Final Handicapped Score  Notes
Michael Madigan 339  400.883  1^{st} Place (turkey)
John Keeney  379  400.583  3^{rd} Place (turkey)
Mark Dennehy  380  400.000  (wine)
Ray Reilly  386  399.142  (ammo)

3.1.11 UCD/TCD Clay Pigeon Colours

FSC Feb 4, 2001

The scoring for this competition was debated at length in the pub afterwards - the agreed-upon result was that it was a good idea to have the clay colours again next year ! (Though special mention to Peter Byrne – missing a perfect score by one clay on his first ever clay pigeon shoot – and Ray Reilly, who has “the most interesting shotgun style ever seen”)

3.1.12 Wilkinstown Open 2

WTSC Jan 13/14, 2001 10m Air Rifle

Name  Class Final Score  Notes
Mark Dennehy C  507  3^{rd} Class C

Also almost won a prize for locking myself out of my car – with the rifle inside ...

3.1.13 Fassaroe 10m Air Rifle Open

UCDRC Feb 17/18, 2001

Name  Class Final Score  Notes
Mark Dennehy  B  528  And a nice Final performance too!
Alan Smyth  C  488  1^{st} Class C
Aisling Broderick C  441
Liam Murray  D  466  2^{nd} Class D
Peter Byrne  D  430

With something on the order of twenty WTSC members entered, we did quite well - if outnumbered!

3.1.14 Rathdrum Open

RRPC Feb 11, 2001 25yd Prone Rifle

Name  Class Final Score  Notes
Mark Dennehy  B  575  4^{th} Overall, 2^{nd} Class B
John Keeney  D  558  2^{nd} Class D
Peter Byrne  D  504
Conor McDermottroe D  523
Michael Madigan  D  525
Liam Murray  D  531

3.1.15 Colours

UCDRC Mar 3, 2001 10m Air Rifle

Colours this year was won in the experienced class by DURC! (And just pipped at the post for the junior class by UCDRC.) But we do get to gloat as this was the first year they didn’t have their “ringers” on their team and were severly trounced as a result.

3.1.16 Rathdrum Spring Special

RRPC April 1, 2001 25yd Prone Rifle

Name  Class Final Score  Notes
Ray Reilly  A  582  3^{rd} Class A
Mark Dennehy  B  578  2^{nd} Class B & 4^{th} Overall
Conor MacDermottroe D  577  1^{st} Class D & 6^{th} Overall – Wow!
John Keeney  C567 1^{st} Class C
Michael Madigan  D  489

Note that this was a 50-point increase by Conor!

3.1.17 National Indoor Prone Rifle Championships

RRPC April 22

Name  Class Final Score  Notes
Mark Dennehy  B  587  National Champion!
(and 1^{st} tun and 1^{st} Class B)
Ray Reilly  A  580
John Keeney  C  574  1^{st} Class C
Conor MacDermottroe C  558
Alan Smyth  C  537
John Keeney  C  558  1^{st} Class C
Peter Byrne  D  553  2^{nd} Class D
Adrian Costigan  D  548
Joe Lawlor  D  547  1^{st} Class D
Liam Murray  D  540
Deirdre Ninabar  D  491
Keith Butler  D  484
Ger Ryan  D  469
Adrian Costigan  D  528

3.2 International

Several internationals this year by Ray : Intershoot, the Hague, World Cup 2000, Milan, and World Cup 2000, Munich.

His Intershoot saga shall never be forgotten - despite having almost all his gear stolen before the competition, he managed to set a personal best ... in Miss Rhona Barry’s shooting clothes ... enough said, really ☺

Unfortunately, his Milan score of 525 proved to be the first symptom of a serious back injury which has put him out of the Air Rifle field for the moment.

4 Internal Competitions

Isle of Man Easter Shoot 2001 – this was my easter holidays and proved to be a disasterous one in Air Rifle, but quite a good one in Prone Rifle. There was a heavy attendance from Irish shooters, and once again, we walked away with a full third of the prizes despite heavy competition.

3.3 Postal Competitions

We entered four postals this year.

UCESSA 10m Air Rifle Postal. Ben is still completing this and doing reasonably well!

UCESSA 25yd Prone Rifle Postal John and I won our divisions of this this year –John only requiring 9 out of 10 cards to do so – and the completed results from last year show John Keeney coming 3^{rd} in his division and me coming 4^{th} in mine. In both years however, some people have started the competition and then not completed their ten cards – next year will see sanctions to prevent this ! ie. you will have to pay your entry fee if you miss more than one card!

NSRA/Eley Postal Competition. This is a very large and prestigous competition and this year we had several entrants, of which four (Ray Reilly, John Keeney, Alan Smyth and Ben Jones) have made it to the second of three rounds.

3.3.1 Trinity Top Ten

The Top Ten this year saw less problems than last year - but almost the same amount of delays. Happily, the scores are now finalised, the prizes made, and have been put in the post to the winners, with the exception of Agnes Graham of Comber, who allready has hers (Agnes may be small in stature, but woe betide those think she’s not scary when she wants to be).

The final Top Ten are :

Name  Club  Score Place
Andrew McFarland EARC 231  1
Gary Duff  Comber 219  2
John Walsh  DRC  218  3
Robert Clarke  EARC 209  4
Richard Farrell  RRPC 209  4
Agnes Graham  Comber 208  6
Tom McKnight  EARC 175  7
Trevor McFarland EARC 171  8
Ronan Power RRPC 163 9
Elizabeth Rodgers EARC 157  10

I’d like to offically thank all the entrants for their patience and for entering – and hopefully they’ll enter again next year when the lessons learnt running it this year are being put into action!

4 Internal Competitions

4.1 Ladders

4.1.1 Prone Rifle

An excellent entry level for Prone this year, with 399 cards shot. The overall best was Ray Reilly with an average of 97.8; which beats his average of last year (97.388) but doesn’t equal his previous year which set the club record of 98.231. Second place was me with 96.657 and John Keeney with 95.629.

The best novices were Adrian Costigan with 92.467, followed closely by Conor McDermottroe and Liam Murray on 91.367 and 91.171 respectivly.

4.1.2 Air Rifle

Air Rifle saw a healthy 331 cards shot this year. The best in the experienced class by a long margin was newcomer Steven Watterson, whose average of 94.000 was well ahead of second place (Me) of 86.200. Third was Miss Aisling Broderick with 80.947 : given that she was studying for scholarships this year, that’s quite respectable.

Best Novice goes to Liam Murray, with 78.333. Well done Liam!

4.1.3 Sporter Rifle

The sporter rifle ladder this year was a lot healthier than last year - this year sees 23 cards as opposed to none last year ! The winner is old hand Frank DeSalle, with an average of 126.25.

4.2 Championships

4.2.1 Prone Rifle

Prone Rifle’s championships were closely fought this year, the winner not being decided until the very last shot fired! The winner, with a score of 98 was Steven Watterson; second on 97 was Mark Dennehy; and third was a tie with both John Keeney and Ray Reilly scoring 96.

4.2.2 Air Rifle

Air Rifle went well this year for Steven, who won with a comfortable 20-point lead, scoring a 96. Second overall and best novice was Adrian Costigan on 76, followed by John Keeney on 75. After reviewing their scores, the rest of us agreed that we’d had enough and a few beers were had with Ben’s chips and salsa to celebrate ☺

4.2.3 Sporter Rifle

Sporter rifle was won by Kenneth Suzuki on 134 : Conor McDemottroe was second with 126 ; and John Keeney came in third on 124.

Last years results weren’t available at the last AGM : they were Willie O’Connor (137); Mark Dennehy (131); and John Keeney (127)

4.2.4 Rapid Fire

Rapid Fire was won by the novices this year, as often happens: Meredith Lulling with 89 and 3 tens won, followed by Peter Byrne on 87 and Me on 86 : though note that with 9 shots, John Keeney managed 6 tens : and a score of 78. Better luck next year John!

Last years results weren’t available at the last AGM : they were Alan Smyth (93); Mark Dennehy (89) and John Keeney(84)

4.3 Other

4.3.1 Standards

Four people shot for standards this year : Michael Madigan, Peter Byrne, John Keeney and myself. Only John was awarded a standard though, and he now receives the badge for the hotshot standard.

5 Club Records

5 Club Records

5.1 .22 calibre Free Rifle

5.1.1 Indoors (25yd) Best Individual Card

Andy Trayler 1993 9× (100, 9 pips shot out)

Highest Ladder Average

Ray Reilly 1999 98.231

Highest Competition Score

Score Ray Reilly 1999 595

(Rathdrum Autumn Open)

Placing Mark Dennehy  2001 Gold Medal, National Championships

Highest Rapid-Fire Score

Derval Tubridy 1993 98

5.1.2 Outdoors Best Individual Card

100yds William Walsh 1987 200(ex200)

Highest Competition Score

50m Prone Ray Reilly 1999 591

(EARC Festival)

50m 3P Simon Kennedy 1995  1084
50m 3P ladies Derval Tubridy 519
Placing Ray Reilly  2000  Gold Medal, National Championships

5.2 .22 calibre Sporting Rifle

5.2.1 Indoors (25yd) Best Individual Card

4P, Sling Keith White  1993 114(ex120)
4P, No Sling Simon Kennedy 1993 107(ex120)

Highest Ladder Average

4P Conor Tighe 1994 107
3P Frank De Salle 2001 126.25

5 Club Records

Highest Competition Score

40shot Prone Simon Kennedy 1994 384(ex400)

5.2.2 Outdoors Highest Competition Score

50m 4P Conor Tighe 1994 160(ex200)
100m Prone Alan Murphy 1990 363(ex400)

5.3 10m Air Rifle

Best Ten-Shot Series

Conor Tighe 1995 99

Highest Ladder Average

Simon Kennedy 1997 94.467

Highest Competition Score

60-shot Ray Reilly 1999 575

FSC 30th Anniversary Shoot

40-shot Derval Tubridy 1997  380

5.4 Fullbore Rifle

Highest Competition Score

200yds M.G.Notley  1963 48(ex50)
300yds M.C.C.Heaton 1962 49(ex50)
500yds P.G.Coulson  ?  48(ex50)
P.Gardiner  ?  48(ex50)
M.G.Notley  ?  48(ex50)
600yds William Walsh ?  48(ex50)
900yds C.Day  ?  45(ex50)
1000yds C.Oakley  ?  45(ex50)

Note that fullbore records date circa 1960 and are spotty at best!

5.5 Miscellaneous Records

3P Air Rifle Andy Traylor 1992  278(ex300)
Triathlon Conor Tighe 1994  298(ex300)
University Top XX Albert Jordan 1998  7 points lost

first Irish shooter to win Individual Gold Medal

Acronyms

Acronyms

D.U.R.C. Dublin University Rifle Club U.C.D.R.C. University College Dublin Rifle Club W.T.S.C. Wilkinstown Target Shooting Club F.S.C. Fassaroe Shooting Club R.R.P.C. Rathdrum Rifle and Pistol Club D.R.C. Dublin Rifle Club E.A.R.C. East Antrim Rifle Club M.R.C. Midlands Rifle Club N.S.R.A. National Small-bore Rifle Association N.T.S.A. National Target Shooting Association N.A.S.R.C. National Association of Sporting Rifle Clubs N.S.A.I. National Silhouette Association of Ireland I.P.C.R.A. Irish Pony Club Rifle Association